# Tractable Gaussian Phase Retrieval with Heavy Tails and Adversarial Corruption with Near-Linear Sample Complexity

### v1

@@ F.1 Guarantees of Algorithm 2 (SampleTopEigenvector) from [22]
$\|\widetilde{\Sigma}_{r}\|_{\mathrm{op}}-\hat{u}^{\top}\widetilde{\Sigma}_{r}\hat{u}\leq O(\gamma)\|\widetilde{\Sigma}_{r}\|_{\mathrm{op}},\quad|\hat{\|x^{*}\|}-\|x^{*}\||\leq\|x^{*}\|/27,\quad\|\Sigma-\widetilde{\Sigma}_{r}\|_{\mathrm{op}}=O(n(K_{4}^{2}+\|x^{*}\|^{4})^{2}/\hat{\tau}^{2})\leq\frac{\sqrt{2}\|x^{*}\|^{4}}{9},$
@@
$dist(\|x^{*}\|u,x^{*})\leq\frac{2\sqrt{2}\|x^{*}\|\|\Sigma-(\widetilde{\Sigma}_{r})\|_{\mathrm{op}}}{12\|x^{*}\|^{4}}\leq\frac{\|x^{*}\|}{27}.$
@@
$5\|x^{*}\|^{4}\|\hat{u}-u\|^{2}\leq O(\gamma)(16\|x^{*}\|^{4}+\sigma^{2})\leq O(\gamma)(16+r_{up}^{2})\|x^{*}\|^{4}\leq 5\|x^{*}\|^{4}/27^{2}.$
So, from above we can conclude that $\|\hat{u}-u\|\leq 1/27.$ So the output $x_{0}$ of $A_{alt}$ satisfies,
$dist\left(x_{0},x^{*}\right)$ $=dist\left(\hat{\|x^{*}\|}\hat{u},x^{*}\right)\leq\|\hat{\|x^{*}\|}\hat{u}-\|x^{*}\|\hat{u}\|+\|\|x^{*}\|\hat{u}-\|x^{*}\|u\|+dist\left(\|x^{*}\|u,x^{*}\right)$
$=|\hat{\|x^{*}\|}-\|x^{*}\||+\|x^{*}\|\|\hat{u}-u\|+dist\left(\|x^{*}\|u,x^{*}\right)\leq\frac{\|x^{*}\|}{9}.$
@@
$5\|x^{*}\|^{4}\|\hat{u}-u\|^{2}=O(\gamma(16\|x^{*}\|^{4}+\sigma^{2}))=O(5\|x^{*}\|^{4}(16+r_{up}^{2}))/(27^{2}(16+r_{up}^{2})).$
So, this implies that $\|\hat{u}-u\|=O(1/27)\leq\frac{C}{27}.$ If the constant $C$ hidden in the big-$O(\cdot)$ notation is less than $1$, then we are done. Otherwise, we set $\gamma=\frac{5}{C^{2}\,27^{2}\,(r_{\mathrm{up}}^{2}+16)},$ and proceed to prove Lemma F.3 using this choice of $\gamma$.
@@ G.1 Local properties of population risk for non-zero mean noise robust phase retrieval
$p$ $=\mathbb{P}(\|X\|\geq\tau)\leq\mathbb{E}[\|X\|^{2}]/\tau^{2}=\mathbb{E}[(\langle b,x^{*}\rangle\langle c,x^{*}\rangle+\zeta)^{2}\|b\|^{2}]/\tau^{2}\leq\mathbb{E}\left[(\langle b,x^{*}\rangle^{2}\langle c,x^{*}\rangle^{2}+\zeta^{2})\left(\sum_{j=1}^{n}b^{2}_{j}\right)\right]/\tau^{2}$
@@
$\|\mathbb{E}[XX^{\top}\mathbb{I}_{\|X\|\geq\tau}]\|_{\mathrm{op}}$ $=\sup_{\|v\|=1}\mathbb{E}[\langle v,X\rangle^{2}\mathbb{I}_{\|X\|\geq\tau}]\leq\sup_{\|v\|=1}\sqrt{p}\sqrt{\mathbb{E}[\langle v,X\rangle^{4}]}\leq\sup_{\|v\|=1}\sqrt{p}\sqrt{\mathbb{E}[8(\zeta^{4}+\langle b,x^{*}\rangle^{4}\langle c,x^{*}\rangle^{4})\langle b,v\rangle^{4}]}$
$\leq\sqrt{p}\sqrt{8(6K_{4}^{4}+2529\|x^{*}\|^{8})}=O(\sqrt{p}\sqrt{(K_{4}^{4}+\|x^{*}\|^{8})})=O(\sqrt{p}(K_{4}^{2}+\|x^{*}\|^{4})).$
@@
$\left|\hat{v}^{2}-(\left\|x^{*}\right\|^{4}+\sigma^{2}/2)\right|$ $=O\left(\sqrt{\operatorname{Var}(v^{2})}\left(\sqrt{2\epsilon}+\sqrt{\frac{\log(4/\delta)}{m_{1}}}\right)\right)=O\left(\sqrt{(r_{up}^{4}+1)\left\|x^{*}\right\|^{8}}\left(\sqrt{2\epsilon}+\sqrt{\frac{\log(4/\delta)}{m_{1}}}\right)\right)$
@@
$\|\hat{u}-u\|^{2}=O\left(\epsilon(r_{up}^{2}+1)+(r_{up}^{2}+1)\sqrt{\epsilon}\right),$ (30)
@@
$=(\lambda_{1}-\lambda_{2})/2\|\hat{u}-u\|^{2}\geq(53/54)\|x^{*}\|^{4}\|\hat{u}-u\|^{2}.$
@@
$\|\hat{u}-u\|^{2}=O\left(\epsilon\left(1+\frac{\sigma^{2}}{\|x^{*}\|^{4}}\right)+73(r_{up}^{2}+1)\sqrt{\epsilon}\right)=O\left(\epsilon(r_{up}^{2}+1)+(r_{up}^{2}+1)\sqrt{\epsilon}\right),$
@@
$|\hat{v}^{2}-(\|x^{*}\|^{4}+\sigma^{2}/2)|\leq\left\|x^{*}\right\|^{4}/54,\quad\|\Sigma-\widetilde{\Sigma}_{r}\|_{\mathrm{op}}\leq\|x^{*}\|^{4}/54,\quad|\lambda_{\hat{u}}-\|\widetilde{\Sigma}_{r}\|_{\mathrm{op}}|=O(\epsilon(r_{up}^{2}+1)+\sqrt{\epsilon}(r_{up}^{2}+1))\|x^{*}\|^{4}=\|x^{*}\|^{4}/27$
@@
$\|\hat{u}-u\|^{2}\leq O\left(\epsilon\left(1+\frac{\sigma^{2}}{\|x^{*}\|^{4}}\right)+(r_{up}^{2}+1)\sqrt{\epsilon}\right)=O\left((r_{up}^{2}+1)\sqrt{\epsilon}\right)\leq(1/27)^{2}$
@@
$dist(\|x^{*}\|u,x^{*})\leq\frac{2\sqrt{2}\|x^{*}\|\|\Sigma-(\widetilde{\Sigma}_{r})\|_{\mathrm{op}}}{2\|x^{*}\|^{4}}\leq\frac{\|x^{*}\|}{27}.$
@@
$dist\left(x_{0},x^{*}\right)$ $=dist\left(\hat{\|x^{*}\|}\hat{u},x^{*}\right)$
$\leq\|\hat{\|x^{*}\|}\hat{u}-\|x^{*}\|\hat{u}\|+\|\|x^{*}\|\hat{u}-\|x^{*}\|u\|+dist\left(\|x^{*}\|u,x^{*}\right)$
$=|\hat{\|x^{*}\|}-\|x^{*}\||+\|x^{*}\|\|\hat{u}-u\|+dist\left(\|x^{*}\|u,x^{*}\right)$
@@
$\mathbb{E}[v^{2}]$ $=\mathbb{E}\left[\left(b^{\top}x^{*}\right)^{2}\left(c^{\top}x^{*}\right)^{2}+\zeta^{2}+2\zeta\left(b^{\top}x^{*}\right)\left(c^{\top}x^{*}\right)\right]=\mathbb{E}\left[\left(b^{\top}x^{*}\right)^{2}\left(c^{\top}x^{*}\right)^{2}+\zeta^{2}\right]=\left\|x^{*}\right\|^{4}+\sigma^{2}/2$
$\operatorname{Var}(v^{2})$ $=\mathbb{E}\left[(\left(b^{\top}x^{*}\right)\left(c^{\top}x^{*}\right)+\zeta)^{4}\right]-(\left\|x^{*}\right\|^{4}+\sigma^{2}/2)^{2}\leq\mathbb{E}\left[8\left(b^{\top}x^{*}\right)^{4}\left(c^{\top}x^{*}\right)^{4}\right]+\mathbb{E}\left[8(\zeta)^{4}\right]-\left\|x^{*}\right\|^{8}$
$\leq 72\left\|x^{*}\right\|^{8}+16K_{4}^{4}-\left\|x^{*}\right\|^{8}\leq 72(r_{up}^{4}+1)\left\|x^{*}\right\|^{8}$
$\mathbb{E}[\upsilon b]$ $=\mathbb{E}[\langle b,x^{*}\rangle\cdot\langle c,x^{*}\rangle b+\zeta b]=\mathbb{E}[\langle b,x^{*}\rangle\cdot\langle c,x^{*}\rangle b]=0,\mathbb{E}[\upsilon c]=\mathbb{E}[\langle b,x^{*}\rangle\cdot\langle c,x^{*}\rangle c+\zeta c]=\mathbb{E}[\langle b,x^{*}\rangle\cdot\langle c,x^{*}\rangle c]=0$
$\mathbb{E}[\upsilon b^{\top}c]$ $=\mathbb{E}[\langle b,x^{*}\rangle\cdot\langle c,x^{*}\rangle b^{\top}c+\zeta b^{\top}c]=\mathbb{E}[\langle b,x^{*}\rangle\cdot\langle c,x^{*}\rangle b^{\top}c]=\mathbb{E}[\sum_{i,j,k=1}^{n}b_{j}x^{*}_{j}c_{i}x^{*}_{i}b_{k}c_{k}]=\sum_{i,j,k=1}^{n}\delta_{\{j=k\}}x^{*}_{j}\delta_{\{i=k\}}x^{*}_{i}$
$=\sum_{k=1}^{n}x^{*2}_{k}=\|x^{*}\|^{2},$
$\operatorname{Var}(vb^{\top}c)$ $=\mathrm{Tr}(\mathbb{E}[\langle b,x^{*}\rangle^{2}bb^{\top}]\mathbb{E}[\langle c,x^{*}\rangle^{2}cc^{\top}])+(\sigma^{2}/2)\mathbb{E}[(b^{\top}c)^{2}]-\|x^{*}\|^{4}=\mathrm{Tr}((\left\|x^{*}\right\|^{2}\mathbb{I}_{n}+2x^{*}x^{*\top})^{2})+(\sigma^{2}/2)n$
$-\|x^{*}\|^{4}=(8+n)\|x^{*}\|^{4}+(\sigma^{2}/2)n-\|x^{*}\|^{4}=n\|x^{*}\|^{4}+n(\sigma^{2}/2)+7\|x^{*}\|^{4}.$
$\operatorname{Cov}[\upsilon b]$ $=\mathbb{E}[\upsilon^{2}bb^{\top}]-\mathbb{E}[\upsilon b]\mathbb{E}[\upsilon b]^{\top}=\mathbb{E}[\upsilon^{2}bb^{\top}]=\mathbb{E}[\langle b,x^{*}\rangle^{2}\cdot\langle c,x^{*}\rangle^{2}bb^{\top}]+\mathbb{E}[\zeta^{2}bb^{\top}]=\|x^{*}\|^{2}\mathbb{E}[\langle b,x^{*}\rangle^{2}bb^{\top}]$
$+\sigma^{2}/2\mathbb{I}_{n}=(\|x^{*}\|^{4}+\sigma^{2}/2)\mathbb{I}_{n}+2\|x^{*}\|^{2}x^{*}x^{*\top}$

### v2

@@ F.1 Guarantees of Algorithm 2 (SampleTopEigenvector) from [24]
$\|\widetilde{\Sigma}_{r}\|_{\mathrm{op}}-\hat{u}^{\top}\widetilde{\Sigma}_{r}\hat{u}\leq O(\gamma)\|\widetilde{\Sigma}_{r}\|_{\mathrm{op}},\quad|\hat{\|x^{*}\|}-\|x^{*}\||\leq\|x^{*}\|/27,$
@@
$\operatorname{dist}(\|x^{*}\|u,x^{*})\leq\frac{2\sqrt{2}\|x^{*}\|\|\Sigma-(\widetilde{\Sigma}_{r})\|_{\mathrm{op}}}{12\|x^{*}\|^{4}}\leq\frac{\|x^{*}\|}{27}\quad\text{and}$
$\operatorname{dist}(-\|x^{*}\|u,x^{*})\leq\frac{2\sqrt{2}\|x^{*}\|\|\Sigma-(\widetilde{\Sigma}_{r})\|_{\mathrm{op}}}{12\|x^{*}\|^{4}}\leq\frac{\|x^{*}\|}{27}.$
@@
$10\|x^{*}\|^{4}(1-\langle\hat{u},u\rangle^{2})\leq O(\gamma)(16\|x^{*}\|^{4}+\sigma^{2})\leq O(\gamma)(16+r_{up}^{2})\|x^{*}\|^{4}\leq 5\|x^{*}\|^{4}/27^{2},$
which further concludes that
$|\langle\hat{u},u\rangle|\geq\sqrt{1-\frac{1}{2\times 27^{2}}}\geq\left(1-\frac{1}{2\times 27^{2}}\right).$
Now,
$\operatorname{dist}(\hat{u},u)$ $=\min\{\|\hat{u}-u\|_{2},\|\hat{u}+u\|_{2}\}$
$=\sqrt{2-2|\langle\hat{u},u\rangle|}=\sqrt{2}\sqrt{1-|\langle\hat{u},u\rangle|}\leq\sqrt{2}\sqrt{1-\left(1-\frac{1}{2\times 27^{2}}\right)}=\frac{1}{27}.$
So, from above we can conclude that $\operatorname{dist}(\hat{u},u)\leq 1/27.$ So the output $x_{0}$ of $A_{alt}$ satisfies,
$\operatorname{dist}\left(x_{0},x^{*}\right)$ $=\operatorname{dist}\left(\hat{\|x^{*}\|}\hat{u},x^{*}\right)\leq\operatorname{dist}(\hat{\|x^{*}\|}\hat{u},\|x^{*}\|\hat{u})+\operatorname{dist}(\|x^{*}\|\hat{u},\|x^{*}\|u)+dist\left(\|x^{*}\|u,x^{*}\right)$
$=\|\hat{\|x^{*}\|}\hat{u}-\|x^{*}\|\hat{u}\|+\operatorname{dist}(\|x^{*}\|\hat{u},\|x^{*}\|u)+dist\left(\|x^{*}\|u,x^{*}\right)$
$=|\hat{\|x^{*}\|}-\|x^{*}\||+\|x^{*}\|\operatorname{dist}(\hat{u},u)+dist\left(\|x^{*}\|u,x^{*}\right)$
@@
$10\|x^{*}\|^{4}(1-\langle\hat{u},u\rangle^{2})\leq O(\gamma)(16\|x^{*}\|^{4}+\sigma^{2})\leq O(\gamma)(16+r_{up}^{2})\|x^{*}\|^{4}\leq 5\|x^{*}\|^{4}/27^{2},$
which again concludes that
$|\langle\hat{u},u\rangle|\geq\sqrt{1-\frac{1}{2\times 27^{2}}}\geq\left(1-\frac{1}{2\times 27^{2}}\right).$
Now,
$\operatorname{dist}(\hat{u},u)$ $=\min\{\|\hat{u}-u\|_{2},\|\hat{u}+u\|_{2}\}$
$=\sqrt{2-2|\langle\hat{u},u\rangle|}=\sqrt{2}\sqrt{1-|\langle\hat{u},u\rangle|}\leq\sqrt{2}\sqrt{1-\left(1-\frac{1}{2\times 27^{2}}\right)}=\frac{1}{27}.$
So, from above we can conclude that $\operatorname{dist}(\hat{u},u)\leq 1/27$. So, this implies that $\|\hat{u}-u\|=O(1/27)\leq\frac{C}{27}.$ If the constant $C$ hidden in the big-$O(\cdot)$ notation is less than $1$, then we are done. Otherwise, we set $\gamma=\frac{5}{C^{2}\,27^{2}\,(r_{\mathrm{up}}^{2}+16)},$ and proceed to prove Lemma F.3 using this choice of $\gamma$. ∎
@@ G.1 Local properties of population risk for non-zero mean noise robust phase retrieval
$p$ $=\mathbb{P}(\|X\|\geq\tau)\leq\mathbb{E}[\|X\|^{2}]/\tau^{2}=\mathbb{E}[(\langle b,x^{*}\rangle\langle c,x^{*}\rangle+\zeta)^{2}\|b\|^{2}]/\tau^{2}=\frac{\mathbb{E}\left[(\langle b,x^{*}\rangle^{2}\langle c,x^{*}\rangle^{2}+\zeta^{2})\left(\sum_{j=1}^{n}b^{2}_{j}\right)\right]}{\tau^{2}}$
@@
$\|\mathbb{E}[XX^{\top}\mathbb{I}_{\|X\|\geq\tau}]\|_{\mathrm{op}}$ $=\sup_{\|v\|=1}\mathbb{E}[\langle v,X\rangle^{2}\mathbb{I}_{\|X\|\geq\tau}]\leq\sup_{\|v\|=1}\sqrt{p}\sqrt{\mathbb{E}[\langle v,X\rangle^{4}]}$
$\leq\sup_{\|v\|=1}\sqrt{p}\sqrt{\mathbb{E}[8(\zeta^{4}+\langle b,x^{*}\rangle^{4}\langle c,x^{*}\rangle^{4})\langle b,v\rangle^{4}]}\leq\sqrt{p}\sqrt{8(6K_{4}^{4}+2529\|x^{*}\|^{8})}$
$=O(\sqrt{p}\sqrt{(K_{4}^{4}+\|x^{*}\|^{8})})=O(\sqrt{p}(K_{4}^{2}+\|x^{*}\|^{4})).$
@@
$\left|\hat{v}^{2}-(\left\|x^{*}\right\|^{4}+\sigma^{2}/2)\right|$ $=O\left(\sqrt{\operatorname{Var}(v^{2})}\left(\sqrt{2\epsilon}+\sqrt{\frac{\log(4/\delta)}{m_{1}}}\right)\right)$
$=O\left(\sqrt{(r_{up}^{4}+1)\left\|x^{*}\right\|^{8}}\left(\sqrt{2\epsilon}+\sqrt{\frac{\log(4/\delta)}{m_{1}}}\right)\right)$
@@
$\operatorname{dist}(\hat{u},u)\leq\sqrt{2-2\sqrt{1-O\left((r_{up}^{2}+1)\sqrt{\epsilon}\right)}},$ (30)
@@
$=106/54(1-\langle\hat{u},u\rangle^{2}).$
So now,
$(1-\langle\hat{u},u\rangle^{2})=O\left(\epsilon\left(1+\frac{\sigma^{2}}{\|x^{*}\|^{4}}\right)+73(r_{up}^{2}+1)\sqrt{\epsilon}\right)=O\left(\epsilon(r_{up}^{2}+1)+(r_{up}^{2}+1)\sqrt{\epsilon}\right)=O\left((r_{up}^{2}+1)\sqrt{\epsilon}\right).$
@@
$|\langle\hat{u},u\rangle|\geq\sqrt{1-O\left((r_{up}^{2}+1)\sqrt{\epsilon}\right)}.$
Now, we can conclude that,
$\operatorname{dist}(\hat{u},u)$ $=\min\{\|\hat{u}-u\|_{2},\|\hat{u}+u\|_{2}\}$
$=\sqrt{2-2|\langle\hat{u},u\rangle|}$
$\leq\sqrt{2-2\sqrt{1-O\left((r_{up}^{2}+1)\sqrt{\epsilon}\right)}}.$
@@
$|\hat{v}^{2}-(\|x^{*}\|^{4}+\sigma^{2}/2)|\leq\left\|x^{*}\right\|^{4}/54,\quad\|\Sigma-\widetilde{\Sigma}_{r}\|_{\mathrm{op}}\leq\|x^{*}\|^{4}/54,$
$|\lambda_{\hat{u}}-\|\widetilde{\Sigma}_{r}\|_{\mathrm{op}}|=O(\epsilon(r_{up}^{2}+1)+\sqrt{\epsilon}(r_{up}^{2}+1))\|x^{*}\|^{4}=\|x^{*}\|^{4}/27.$
@@
$\operatorname{dist}(\hat{u},u)\leq\sqrt{2-2\sqrt{1-O\left((r_{up}^{2}+1)\sqrt{\epsilon}\right)}}\leq(1/27),$
@@
$\operatorname{dist}(\|x^{*}\|u,x^{*})\leq\frac{2\sqrt{2}\|x^{*}\|\|\Sigma-(\widetilde{\Sigma}_{r})\|_{\mathrm{op}}}{2\|x^{*}\|^{4}}\leq\frac{\|x^{*}\|}{27}\quad\text{and}$
$\operatorname{dist}(-\|x^{*}\|u,x^{*})\leq\frac{2\sqrt{2}\|x^{*}\|\|\Sigma-(\widetilde{\Sigma}_{r})\|_{\mathrm{op}}}{2\|x^{*}\|^{4}}\leq\frac{\|x^{*}\|}{27}.$
@@
$\operatorname{dist}\left(x_{0},x^{*}\right)$ $=\operatorname{dist}\left(\hat{\|x^{*}\|}\hat{u},x^{*}\right)\leq\operatorname{dist}(\hat{\|x^{*}\|}\hat{u},\|x^{*}\|\hat{u})+\operatorname{dist}(\|x^{*}\|\hat{u},\|x^{*}\|u)+dist\left(\|x^{*}\|u,x^{*}\right)$
$=\|\hat{\|x^{*}\|}\hat{u}-\|x^{*}\|\hat{u}\|+\operatorname{dist}(\|x^{*}\|\hat{u},\|x^{*}\|u)+dist\left(\|x^{*}\|u,x^{*}\right)$
$=|\hat{\|x^{*}\|}-\|x^{*}\||+\|x^{*}\|\operatorname{dist}(\hat{u},u)+dist\left(\|x^{*}\|u,x^{*}\right)$
@@
$\mathbb{E}[v^{2}]$ $=\mathbb{E}\left[\left(b^{\top}x^{*}\right)^{2}\left(c^{\top}x^{*}\right)^{2}+\zeta^{2}+2\zeta\left(b^{\top}x^{*}\right)\left(c^{\top}x^{*}\right)\right]=\mathbb{E}\left[\left(b^{\top}x^{*}\right)^{2}\left(c^{\top}x^{*}\right)^{2}+\zeta^{2}\right]$
$=\left\|x^{*}\right\|^{4}+\sigma^{2}/2$
$\operatorname{Var}(v^{2})$ $=\mathbb{E}\left[(\left(b^{\top}x^{*}\right)\left(c^{\top}x^{*}\right)+\zeta)^{4}\right]-(\left\|x^{*}\right\|^{4}+\sigma^{2}/2)^{2}\leq\mathbb{E}\left[8\left(b^{\top}x^{*}\right)^{4}\left(c^{\top}x^{*}\right)^{4}\right]+\mathbb{E}\left[8(\zeta)^{4}\right]$
$-\left\|x^{*}\right\|^{8}\leq 72\left\|x^{*}\right\|^{8}+16K_{4}^{4}-\left\|x^{*}\right\|^{8}\leq 72(r_{up}^{4}+1)\left\|x^{*}\right\|^{8}$
$\mathbb{E}[\upsilon b]$ $=\mathbb{E}[\langle b,x^{*}\rangle\cdot\langle c,x^{*}\rangle b+\zeta b]=\mathbb{E}[\langle b,x^{*}\rangle\cdot\langle c,x^{*}\rangle b]=0,\mathbb{E}[\upsilon c]=\mathbb{E}[\langle b,x^{*}\rangle\cdot\langle c,x^{*}\rangle c+\zeta c]$
$=\mathbb{E}[\langle b,x^{*}\rangle\cdot\langle c,x^{*}\rangle c]=0$
$\mathbb{E}[\upsilon b^{\top}c]$ $=\mathbb{E}[\langle b,x^{*}\rangle\cdot\langle c,x^{*}\rangle b^{\top}c+\zeta b^{\top}c]=\mathbb{E}[\langle b,x^{*}\rangle\cdot\langle c,x^{*}\rangle b^{\top}c]=\mathbb{E}\left[\sum_{i,j,k=1}^{n}b_{j}x^{*}_{j}c_{i}x^{*}_{i}b_{k}c_{k}\right]$
$=\sum_{i,j,k=1}^{n}\delta_{\{j=k\}}x^{*}_{j}\delta_{\{i=k\}}x^{*}_{i}=\sum_{k=1}^{n}x^{*2}_{k}=\|x^{*}\|^{2},$
$\operatorname{Var}(vb^{\top}c)$ $=\mathrm{Tr}(\mathbb{E}[\langle b,x^{*}\rangle^{2}bb^{\top}]\mathbb{E}[\langle c,x^{*}\rangle^{2}cc^{\top}])+(\sigma^{2}/2)\mathbb{E}[(b^{\top}c)^{2}]-\|x^{*}\|^{4}=\mathrm{Tr}((\left\|x^{*}\right\|^{2}\mathbb{I}_{n}+2x^{*}x^{*\top})^{2})$
$+(\sigma^{2}/2)n-\|x^{*}\|^{4}=(8+n)\|x^{*}\|^{4}+(\sigma^{2}/2)n-\|x^{*}\|^{4}=n\|x^{*}\|^{4}+n(\sigma^{2}/2)+7\|x^{*}\|^{4}.$
$\operatorname{Cov}[\upsilon b]$ $=\mathbb{E}[\upsilon^{2}bb^{\top}]-\mathbb{E}[\upsilon b]\mathbb{E}[\upsilon b]^{\top}=\mathbb{E}[\upsilon^{2}bb^{\top}]=\mathbb{E}[\langle b,x^{*}\rangle^{2}\cdot\langle c,x^{*}\rangle^{2}bb^{\top}]+\mathbb{E}[\zeta^{2}bb^{\top}]$
$=\|x^{*}\|^{2}\mathbb{E}[\langle b,x^{*}\rangle^{2}bb^{\top}]+\sigma^{2}/2\mathbb{I}_{n}=(\|x^{*}\|^{4}+\sigma^{2}/2)\mathbb{I}_{n}+2\|x^{*}\|^{2}x^{*}x^{*\top}$